# The Changes of the Northern Hadley Cell Strength in Reanalyses and Radiosonde Observations

### v1

@@ 4 Conclusions and Outlook
Focusing on the NHC, our assessment of mean meridional winds reveals that the NHC is generally stronger in observations than in reanalyses, primarily due to stronger upper-tropospheric poleward flow. Additional investigation of ERA5 first-guess departures and analysis increments indicates that the ERA5 forecast model systematically underestimates the NHC strength. Data assimilation corrects it by imposing meridional wind increments that strengthen the circulation, confirming that ERA5 has a negative bias in the NHC strength.
Consistent with previous studies of (Nguyen et al., 2013; Chemke and Polvani, 2019; Zaplotnik et al., 2022; Pikovnik et al., 2022), reanalyses suggest a significant strengthening of the NHC, driven mainly by an intensification of upper-tropospheric poleward flow. Conversely, lower-tropospheric equatorward flow is weakening in the reanalyses, in line with the reported decline in the meridional gradient of mean sea level pressure (Chemke and Yuval, 2023). Contrary to the findings in this study and contrary to climate model projections, our research based on radiosonde observations indicates a neutral trend in NHC strength over the 1980-2022 period and an insignificant strengthening over 1980-2015. Furthermore, we show that the analysis increments do not contribute to the strengthened NHC trend in ERA5.
Our findings (1) raise critical questions about the reliability of reanalyses to estimate long-term HC strength trends, particularly in the context of climate change; (2) suggest that the reanalyses based on surface observations are insufficient to properly constrain the full vertical extent of atmospheric circulation; (3) underscore the essential role of conventional upper-air observations in verifying global circulation changes, particularly over remote ocean areas; (4) raise concerns about the reliability of climate model projections, which indicate a weakening of the NHC.
@@ Acknowledgements
(https://psl.noaa.gov/data/gridded/data.20thC_ReanV3.html). The MERRA-2 reanalysis datasets are freely available at https://disc.gsfc.nasa.gov/ web page, managed by the NASA Goddard Earth Sciences (GES) Data and Information Services Center (DISC).
Code availability. The introduced analysis was performed using Python. The codes used in this study will be open-access and available through a designated Github profile or open Zenodo repository.

### v2

@@ 4 Conclusions and Outlook
Focusing on the NHC, our assessment of mean meridional winds reveals that the NHC is generally stronger in observations than in reanalyses, primarily due to stronger upper-tropospheric poleward flow. Additional investigation of ERA5 first-guess departures and analysis increments indicates that the ERA5 forecast model systematically underestimates the NHC strength. Data assimilation corrects it by imposing meridional wind increments that strengthen the circulation, confirming that ERA5 has a negative bias in the NHC strength. Our results are consistent with Mayer et al. (2025), who showed that the data assimilation in ERA5 enhances the surface meridional winds over the Pacific. Using a completely different approach, Vonich and Hakim (2025) showed that optimal initial-condition perturbations for improving weather predictability involve large-scale corrections to ERA5, indicative of a stronger Hadley circulation, further suggesting that ERA5 underestimates the strength of the Hadley circulation.
Consistent with previous studies of (Nguyen et al., 2013; Chemke and Polvani, 2019; Zaplotnik et al., 2022; Pikovnik et al., 2022), reanalyses suggest a significant strengthening of the NHC in the recent decades, driven mainly by an intensification of upper-tropospheric poleward flow. Conversely, lower-tropospheric equatorward flow is weakening in the reanalyses, in line with the reported decline in the meridional gradient of mean sea level pressure (Chemke and Yuval, 2023). Contrary to the findings in this study and contrary to climate model projections, our research based on radiosonde observations indicates a neutral trend in NHC strength over the 1980-2022 period and an insignificant strengthening over 1980-2015. Furthermore, we show that the analysis increments do not contribute to the strengthened NHC trend in ERA5.
Our findings (1) raise critical questions about the reliability of reanalyses to estimate long-term HC strength trends, particularly in the context of climate change; (2) suggest that the reanalyses based on surface observations are insufficient to constrain the full vertical extent of atmospheric circulation properly; (3) underscore the essential role of conventional upper-air observations in verifying global circulation changes, particularly over remote ocean areas; (4) raise concerns about the reliability of climate model projections, which indicate a weakening of the NHC.
@@ Acknowledgements
(https://psl.noaa.gov/data/gridded/data.20thC_ReanV3.html). The MERRA-2 reanalysis datasets are freely available at the https://disc.gsfc.nasa.gov/ web page, managed by the NASA Goddard Earth Sciences (GES) Data and Information Services Center (DISC).
Code availability. The introduced analysis was performed using Python. The codes used in this study will be open-access and available through a designated GitHub profile or an open Zenodo repository.